OpenStudy (anonymous):

Subtract. (4x + 17xy – 9y) – (5x – 8xy – y) (Points : 15)

OpenStudy (anonymous):

(4x + 17xy – 9y) – (5x – 8xy – y) =4x + 17xy – 9y – 5x + 8xy + y =-x+25xy-8y